GOING FURTHER 1. Based on the investigation, why do you think heat and alcohols are used to disinfecet medical equipment? 2. Sometimes when people accidentally ingest heavy metal ion like mercury they are asked to drink some milk. Why do you think this is? 3. How does a change in pH affect the secondary and tertiary structures of a protein? 4. Calculate the % Casein in your milk sample. (Show work) Suppose you filled one test tube with alanine and a second test tube with casein, but forgot to label the test tubes. What test could you run on the samples to confirm which test tube contained the alanine? Explain. 5. 6. In the colorimetric test, did your results agree with what you were supposed to see?
Answer #1

1. Alcohol is used because this is poisonous, so when we use it all microbes presents in the medical equipment die with the contact. When we usea heat, this desnature the proteins in the microbes cells, causing damage of its and finally the organism dies.

2. Heavy metal ions causes damage at the protein sites on aour body, when you drink milk after ingest accidentally the heavy metal, the heavy metal ions will "attack" on the milk proteins rather than the our body proteins. This produces that person vomit up the milk and the heavy metal ingested.

3. A change in pH affects the aqueous envelope proteins and affect the electric charge of the acidic and basic groups in the side chains of amino acids . This alteration of the surface charge of the protein eliminates electrostatic interactions that stabilize the tertiary structure and often causes precipitation.

4. Calculated mass of milk = 85.37 - 64.41 = 20.96 g

Calculated mass of casein = 23.71 - 22.91 = 0.8 g

% Casein = (mass os casein/mass of milk)*100 = (0.8g/20.96g)*100

% Casein = 3.82 %

5. To identify the alanina we could use a Ninhydrin reaction because in this reaction The alpha- amino group of amino acids complexed colored with ninhydrin : violet bluish in most amino acids whose amino group is primary, yellow for proline and hydroxyproline and coffee for asparagine having an amido group in the side chain. This
reaction also identifies the free alpha-amino groups in peptides and proteins.